OpenStudy (anonymous):

State the value of the discriminant and then describe the nature of the solutions. -14x^2+5x-8=0 What is the valuel of the discriminant? which is correct: The equation has one real solution? has two imaginary solutions? has two real solutions?

OpenStudy (amistre64):

discriminate is just: \(middle^2-4(first)(last)\)

OpenStudy (amistre64):

5^2 -4(-14)(-8)

OpenStudy (amistre64):

i see it as imaginary solutons cause the answer is gonna be negative

OpenStudy (anonymous):

k

OpenStudy (amistre64):

make sense?

OpenStudy (anonymous):

not really, i thought it would have been a positive since there are two negatives in the equation

OpenStudy (amistre64):

5^2 = 25 -4(-8) = 32 32(-14) = -320-120-8 = -448 25-448 = - something lol

OpenStudy (amistre64):

to discriminate means to make a choice based upon some criteria; to discrimination based on skin color is an obvious correlation. the discriminate of a quadratic equation chooses how many roots it has.

OpenStudy (amistre64):

if the discriminate is +; we will get 2 roots if it = 0; there is only 1 root if it is -; ony imaginary roots exists

OpenStudy (anonymous):

ok i think I see